OpenStudy (anonymous):

Helena made trail mix and wants to list the ingredients in order of increasing weight. She used one and five eighths lb of mixed nuts, five eighths lb of chocolate covered raisins, and one and two eighthslb of dried mixed fruit. If she decides

OpenStudy (anonymous):

to include three eighths lb of sunflower seeds, which list shows the order of ingredients from least weight to greatest weight

OpenStudy (anonymous):

raisins, seeds, fruit, nuts seeds, raisins, fruit, nuts raisins, seeds, nuts, fruit seeds, raisins, nuts, fruit

OpenStudy (anonymous):

um lets see... convert them all to improper fractions seeds:3/8 raisins:5/8 fruit:1 2/8 = 10/8 nuts:1 5/8 = 13/8

OpenStudy (anonymous):

least to greatest would be : seeds, raisins, fruit, then nuts

OpenStudy (anonymous):

so the second one

OpenStudy (anonymous):

@Juliette_Rose I have another question would you like to help me

OpenStudy (anonymous):

did u get my answer though? and sure

OpenStudy (anonymous):

yes Which of the following is not true? 2 and 5 are the only prime factors of twenty. The prime factorization of forty-six is 2 times twenty-three. 3 is the only prime factor of twenty-seven. The prime factorization of twenty-four is 2 squared times 3 squared.

OpenStudy (anonymous):

@Juliette_Rose

OpenStudy (anonymous):

Derek distributed twenty pieces of apple to his friends. If each apple slice was one-eighth of an apple, how many apples did he distribute to his friends? three and one-fourth two and three-fourths three and one-eighth two and one-half

OpenStudy (anonymous):

@Juliette_Rose

OpenStudy (anonymous):

for the 2nd one about apples: 1/8 times 20=20/8= 2 and a 1/2

OpenStudy (anonymous):

so the second one is ....... im confused

OpenStudy (anonymous):

How'd I lose you?

OpenStudy (anonymous):

with the answer for second @Juliette_Rose

OpenStudy (anonymous):

@Juliette_Rose

OpenStudy (anonymous):

so the answer for the apple thing is 1/2/2

OpenStudy (anonymous):

no 1 1/2as in one and one-half

OpenStudy (anonymous):

What is the Least Common Denominator (LCD) of nine-sixteenths and nineteen twenty-fourths? forty-eight thirty-two sixteen 4 @Juliette_Rose I also need help with alot

OpenStudy (anonymous):

@Juliette_Rose

OpenStudy (anonymous):

@Juliette_Rose

OpenStudy (anonymous):

@Juliette_Rose

OpenStudy (anonymous):

@Juliette_Rose

OpenStudy (anonymous):

@Juliette_Rose

OpenStudy (anonymous):

@Juliette_Rose

OpenStudy (anonymous):

honestly i havent done most of these since the 4th grade and i still have a ton of my own problems to do... Im really sorry

OpenStudy (anonymous):

ohhhh man I really wish u could've just answered this last question :(

OpenStudy (anonymous):

@Juliette_Rose

OpenStudy (anonymous):

i havent even made a dent in my own he ^_^" sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@help123please. @bd3ll

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

i need help @help123please.